• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Z. Chai

    Log entry time 07:55:18 on September 26,1999

    Entry number 27201

    keyword=Owl Shift Summary

    00:00---Still waiting for beam.
    DAQ is taking data for Run 1838 without beam in the Hall A.

    00:31---MCC told us they still need one more hour to send beam to Hall A.

    01:02---Stopped Run 1838 ( .4M evts)
    MCC is still tuning the beam.

    01:23---MCC told us they can send beam to Hall A now.
    Ask them to send 5 microA to check the raster first.

    01:30---Beam 3.8 microA.
    Start Run 1839 for spot check.(LD2 15cm target)
    Beam is not stable yet.

    01:36---Spot size 1.5/2.5mm
    The beam is tripping frequently.

    01:47---The beam seems stable now.
    Ask MCC to send us the maximum beam.
    I_beam = 17microA.

    01:48---Start Run 1840(LD2 15cm, radiator in 6%, P_e=2.60880Gev/c,
    P_h=1.93660Gev/c, Theta_e =27.505, Theta_h=51.170, All carbon door are closed 1.5+3+6+9)

    02:16---Replayed Run 1840. Everything looks fine.
    Beam is pretty stable now. ~17 microA.

    02:42---Beam dropped to 16.3microA.

    02:51---Beam tripped for half a minute.

    02:53---Stop Run 1840 (903K evts, .06C)

    02:54---Start Run 1841
    Beam ~16 microA

    03:58---Stop Run 1841(869K evts, .06C)
    The total charge from Run 1831 to Runb 1841 is 0.471C for (gamma d)
    Move the radiator out

    04:01---Start Run 1842(LD2 15 cm, radiator out)

    05:01---Stop Run 1842(209K evts, .05C)

    05:02---Start Run 1843

    06:01---Stop Run 1843(200K evts, .05C)

    06:02---Start Run 1844
    Beam 14.3 microA

    07:00---Stop Run 1844(187K evts,.047C)

    07:01---Start Run 1845

    08:00---Continue data taking Run 1845